# Shannon entropy of optimized proton-neutron pair condensates

### v1

@@ IV Summary and Outlook
In summary, we have implemented angular momentum projection after variation of an explicit general proton-neutron pair condensate (PN-PC), with good particle numbers, under shell model effective interactions. With one optimized proton-neutron pair condensate as the reference state, we benchmark PN-PC against shell model results, the Projected Hartree-Fock method, and experimental data when available. Qualitative agreements are achieved on yrast excitation energies and B(E2) values for a series of even-even $N=Z$ nuclei in the $1s0d$ and $1p0f$ major shell, though the PN-PC one-reference ground state energies are around 1-2 MeV above that of the shell model. In the $1p0f_{5/2}0g_{9/2}$ major shell, one unique proton-neutron pair condensate does not reproduce the vibrational-type yrast spectrum of 88Ru, 92Pd and 96Cd, but qualitatively reproduces the B(E2) values among yrast states. A dual description is observed in 92Pd, i.e. the optimized T=0 proton-neutron condensate and the optimized T=1 proton-neutron condensate generate similar spectrum and B(E2) values. Analysis of Shannon correlation entropy reveals that, like-nucleon pair condensates in semi-magic nuclei show an entangled phase, while optimized proton-neutron pair condensates in $N=Z$ nuclei are not far from a mean-field Slater determinant, as a result of the complicity of both T=0 and T=1 interactions.
@@ Acknowledgements.
We would like to thank discussions with Xin Guan on entropy of pairing configurations. Yi Lu acknowledges support by the Natural Science Foundation of Shandong Province, China (ZR2022MA050), the National Natural Science Foundation of China (11705100, 12175115). Yang Lei is grateful for the financial support of the Sichuan Science and Technology Program (Grant No. 2019JDRC0017), and the Doctoral Program of Southwest University of Science and Technology (Grant No. 18zx7147). This material is also based upon work (Calvin W. Johnson) supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under Award Number DE-FG02-03ER41272.

### v2

@@ IV Summary and Outlook
In summary, we have implemented angular momentum projection after variation of an explicit general proton-neutron pair condensate (PN-PC), with good particle numbers, under shell model effective interactions. With one optimized proton-neutron pair condensate as the reference state, we benchmark PN-PC against shell model results, the Projected Hartree-Fock method, and experimental data when available. Qualitative agreements are achieved on yrast excitation energies and B(E2) values for a series of even-even $N=Z$ nuclei in the $1s0d$ and $1p0f$ major shell, though the PN-PC one-reference ground state energies are around 1-2 MeV above that of the shell model. In the $1p0f_{5/2}0g_{9/2}$ major shell, one unique proton-neutron pair condensate does not reproduce the vibrational-type yrast spectrum of 88Ru, 92Pd and 96Cd, but qualitatively reproduces the B(E2) values among yrast states. A dual description is observed in 92Pd, i.e. the optimized T=0 proton-neutron condensate and the optimized T=1 proton-neutron condensate generate similar spectrum and B(E2) values. Analysis of Shannon correlation entropy reveals that, like-nucleon pair condensates in semi-magic nuclei show an entangled phase, while optimized proton-neutron pair condensates in $N=Z$ nuclei from 18F to 98In are not far from a mean-field Slater determinant, as a result of the complicity of both T=0 and T=1 interactions.
@@ Acknowledgements.
We would like to thank discussions with Xin Guan on Shannon entropy of pairing configurations. Yi Lu acknowledges support by the Natural Science Foundation of Shandong Province, China (ZR2022MA050), the National Natural Science Foundation of China (11705100, 12175115). Yang Lei is grateful for the financial support of the Sichuan Science and Technology Program (Grant No. 2019JDRC0017), and the Doctoral Program of Southwest University of Science and Technology (Grant No. 18zx7147). This material is also based upon work (Calvin W. Johnson) supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under Award Number DE-FG02-03ER41272. Guan-Jian Fu acknowledges the National Natural Science Foundation of China (12075169, 12322506), the Fundamental Research Funds for Central Universities (22120240207).